OpenStudy (anonymous):

x^2+10x+21 factor completely

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} x^2&+10x&+21\\ &\uparrow &\uparrow \\ &7+3&7\cdot 3 \end{array}\)

OpenStudy (anonymous):

Im still not getting it. How did you got that answer?. can you explain please?

OpenStudy (jdoe0001):

have you covered this factoring quadratics yet? should be pretty obvious

OpenStudy (anonymous):

Nope not yet, that's the next assignment I have to work on.

OpenStudy (jdoe0001):

you'd end up with 2 binomials, (x + something1)(x + something2) an the "something1" and "something2" should sum up to the middle term and multiply to produce the last term

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} x^2&+10x&+21\\ &\uparrow &\uparrow \\ &7+3&7\cdot 3 \end{array}\bf \implies (x+7)(x+3)\)

OpenStudy (jdoe0001):

you can try multiplying the binomials, you'd end up with the original polynomial

OpenStudy (anonymous):

ohhhhh alright I'm getting it now and you use the FOIL method right?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

Alright, I got it thanks you so much :)

OpenStudy (jdoe0001):

yw